OpenStudy (lilcathyx):

Im gonna be honest i need help with 3 questions. that i seriously dont know anything about.. im close to giving up on this...

OpenStudy (lilcathyx):

1. The theme park company is building a scale model of the killer whale stadium main how tank for an investor's presentation. Each dimension will be made 6 times smaller to accommodate the mock-up in the presentation room. How many times smaller than the actual volume is the volume of the mock-up? 2. Using the same information from #5, what percent of change occurred from the actual tank to the mock-up of the tank? 3. If you were to take a cross section parallel to the base of one of the holding tanks, how would you describe the shape?

OpenStudy (lilcathyx):

@SilenTempest @jim_thompson5910 ive been suck on these questions for dayss. :/ -__-

jimthompson5910 (jim_thompson5910):

think of a cube imagine it is 6 by 6 by 6 what is the volume of this cube?

OpenStudy (lilcathyx):

216 ?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

now divide each dimension by 6 6 by 6 by 6 ----> 1 by 1 by 1

jimthompson5910 (jim_thompson5910):

the new volume is??

OpenStudy (lilcathyx):

1 ?

jimthompson5910 (jim_thompson5910):

so the old volume is 216, the new is 1 you divide the old volume by 216 to get the new volume you'll find that this trick works for any type of cube

jimthompson5910 (jim_thompson5910):

let's say in general, the side lengths are s so the volume of the cube is s^3 if you divide each dimension by k, then the new volume of the smaller cube will be \[\Large \frac{s^3}{k^3}\]

jimthompson5910 (jim_thompson5910):

and this idea applies to any 3D shapes (not just cubes)

OpenStudy (lilcathyx):

ok this is all for 1 ? im ngl but im a bit lost, just a bit.

jimthompson5910 (jim_thompson5910):

yeah this is going to help you with #1

jimthompson5910 (jim_thompson5910):

do another example let's say the cube now has sides that are 10 inches long (so s = 10) the volume of this cube is ______ then divide each dimension by 5 (so k = 5) the volume of the new smaller cube is ______

jimthompson5910 (jim_thompson5910):

compare the new volume to the expression (s^3)/(k^3) when s = 10 and k = 5

OpenStudy (lilcathyx):

1000 is the cubes volume (s= 10)

OpenStudy (lilcathyx):

k = 125

OpenStudy (lilcathyx):

1000/125 = 8 right? @jim_thompson5910

OpenStudy (lilcathyx):

the new volume is 8 ?

OpenStudy (lilcathyx):

@jim_thompson5910 ^.^ :o

jimthompson5910 (jim_thompson5910):

that is correct also notice how 10/5 = 2 and 2^3 = 8 so the new volume is 1/8 of the old volume

jimthompson5910 (jim_thompson5910):

that's for the example though

OpenStudy (lilcathyx):

yesssssss. :o lets see if i understand so in Q 1 : "s" S =6 right ? and K = nothing. just 6^3 ?

jimthompson5910 (jim_thompson5910):

s is unknown for #1 k = 6 because we're dividing each dimension by 6

jimthompson5910 (jim_thompson5910):

so old volume = s^3 new volume = (s^3)/(k^3) = (1/k^3)*s^3 notice how I pulled out the 1/k^3

jimthompson5910 (jim_thompson5910):

new volume = (1/k^3)*s^3 new volume = (1/6^3)*s^3 new volume = (1/216)*s^3 so we see that the new volume is 1/216 of the old volume (basically it's 216 times smaller)

OpenStudy (lilcathyx):

so its 216 times smaller ? lmfaoo sorry i was copying that in my notebook

jimthompson5910 (jim_thompson5910):

yes it is

jimthompson5910 (jim_thompson5910):

if we divided everything by 4 instead, it would be 4^3 = 64 times smaller

OpenStudy (lilcathyx):

is it still right though ?

jimthompson5910 (jim_thompson5910):

yeah I was just bringing up another example

OpenStudy (lilcathyx):

ohhh i see.

OpenStudy (lilcathyx):

for #2 all i have to do is turn it into percentage and subtract :?

jimthompson5910 (jim_thompson5910):

what was the volume from #5 ?

OpenStudy (lilcathyx):

thats actually #1 i forgot to change it. its not #5

jimthompson5910 (jim_thompson5910):

ok let's say we go from old volume = 216 new volume = 1 what is the percent change?

OpenStudy (lilcathyx):

99.537 --> 99.5 ?

jimthompson5910 (jim_thompson5910):

that's correct

OpenStudy (lilcathyx):

thats all for # 2 ? Whatt ?

jimthompson5910 (jim_thompson5910):

yeah we have roughly a 99.5% reduction

OpenStudy (lilcathyx):

im finally understanding now ! :D

jimthompson5910 (jim_thompson5910):

I'm glad you are

OpenStudy (lilcathyx):

For #3 and i promise i wont bother anymore, would it be a cube or a triangle or a sphere?

jimthompson5910 (jim_thompson5910):

what 3D shape are the holding tanks again?

OpenStudy (lilcathyx):

sphere?

jimthompson5910 (jim_thompson5910):

they mention a base though, so that's a bit odd

OpenStudy (lilcathyx):

yup..

jimthompson5910 (jim_thompson5910):

well if it is a sphere, then yes, the cross section will be a circle

jimthompson5910 (jim_thompson5910):

imagine the shadow of the sphere

jimthompson5910 (jim_thompson5910):

but the base thing is throwing me off

OpenStudy (lilcathyx):

me too...

OpenStudy (lilcathyx):

thats why i thought of a sphere